1. A water surface wave with a frequency of 0.379 Hz has a wavelength of 16.7 m. What is its wave speed?

2. A water surface wave has a wave speed of 16.3 m/s and a frequency of 0.110 Hz. What is its wavelength?
Answer #1

1. v = wavelength * frequency = 0.379 Hz * 16.7 m = 6.329 m/s

2. wavelength = v/frequency = 16.3 m/s /0.110 Hz = 148.18 m
Answer #2

u10l2e2.gif

speed=
Answer #3

Here is what I solved before, please modify the figures as per your question. Please let me know if you have further questions. Ifthis helps then kindly rate 5-stars.

A water surface wave with a frequency of 0.471 Hz has a wavelength of 17.8 m. What is its wave speed?

answer

v=f?

=0.471*17.8 [f=frequency]

   =8.38m/s

A water surface wave has a wave speed of 16.4 m/s and a frequency of 0.060 Hz. What is its wavelength   

answer

wave length = velocity/frequency = 16.4/0.06 = 273.333 m   
Answer #4

Part 1)

Apply v = f(wavelength)

v = (.379)(16.7)

v = 6.33 m/s

Part 2)

Same formula

16.3 = .110(wavelength)

wavelength = 148.2 m
Answer #5

velocity of a sound wave ( wave speed ) = wave length * frequency = 16.7 *0.379

= 6.18 m/s

wave length = wave speed / frequency

= 16.3/0.11

   = 148.18 m